Fill in the table below for the following zero-coupon bonds, all of which have par values of $1,000. Use semi-annual periods. (Do not round intermediate calculations. Round your answers to 2 decimal places.).

my hw is due in 45 minutes please help!!

Price Maturity (years) Yield to Maturity
$480 20 %
$580 20 %
$580 10 %
10 10.80 %
10 7.20 %
$480 8.80 %
Answer #1

1
=((1000/480)^(1/(2*20))-1)*2=3.70372218142481%

2
=((1000/580)^(1/(2*20))-1)*2=2.74226583081112%

3
=((1000/580)^(1/(2*10))-1)*2=5.5221317710564%

4
=1000/(1+10.8%/2)^(2*10)=349.291311195797

5
=1000/(1+7.2%/2)^(2*10)=492.952293807799

6
=1/2*LOG(1000/480)/LOG(1+8.8%/2)=8.52273429477606